Session: Feb 1999 (session 1)
Wavelength: 6 cm


To give feedback on this experiment, click here

Experiment Feedback

Medicina (Orfei at Mon Feb 8 16:14:42 MET 1999) Minor failures
Observation ok. The right channel has a tsys of about 300K due to some problem on crio preamlifier. Anyway, from preliminary checks DA193 showed detectable.

Torun (K.Borkowski at Mon Feb 8 16:21:44 MET 1999) Minor failures
No 1pps to BBC 03. No Tsys measurements.

Westerbork (Tony Foley at Mon Feb 8 16:28:33 MET 1999) Minor failures
Using the apex phasecal injection I found first 4 VCs had correct polarization (odd=LCP in VC01,VC03) final 10 VCs had SWAPPED polarization (odd=RCP in VC05-VC13) Estimated SEFD about 220Jy, MUCH worse than I expected. Tape started late so please ignore pass 1 and most of pass 2. Passes 3 and 4 look reasonable.

Jodrell (Alastair Gunn at Mon Feb 8 16:54:54 MET 1999) Success
Snow on antenna dish may have reduced sensitivity. No other known problems.

Cambridge (Alastair Gunn at Mon Feb 8 16:55:36 MET 1999) Failure
Cambridge did not observe FT002 due to failure of recorder heads.

Effelsberg (dgraham at Tue Feb 9 11:43:53 MET 1999) Success
Seems OK, MK4 clock has an unusual offset , 3musec earlier than VLBA. Zenith Tsys 30K, Tant for DA193 about 8K.

Noto (Trigilio at Tue Feb 9 16:47:56 MET 1999) Success
No problem during the observations. Strong wind.

Noto (Trigilio at Tue Feb 9 16:48:03 MET 1999) Success
No problem during the observations. Strong wind.

MPIfR (dgraham at Thu Feb 11 15:29:52 MET 1999) Success
Correlator report so far: tapes received Effelsberg/Onsala/Jodrell/Torun, fringes with correct polarization on all. Further details by Email.

Shanghai (Huang Xinyong at Sun Feb 21 02:56:08 MET 1999) Success
No problems during observating.

Shanghai (Huang Xinyong at Mon Feb 22 10:53:59 MET 1999) Success
It is cloudy.

Noto (C.Trigilio at Fri Feb 26 11:31:31 MET 1999) Minor failures
BBC12 not locked to the right frequency.